OpenStudy (anonymous):

The Discount Hair Center has a large container of combs. Each purchase qualifies a customer to receive a free comb. The container contains 4 blue combs, 2 purple combs, 8 black combs, 3 white combs, 6 yellow combs, and 5 pink combs. What is the probability that the next customer will pull out a blue comb?

OpenStudy (anonymous):

4/total possibilities =4/(4+2+8+3+6+5) =4/28 =1/7